1. What is the ionic strength of a solution composed of 0.M NaCl and 0.2M Na2SO4?


1
Expert's answer
2022-05-12T02:49:02-0400

"NaCl \rightarrow Na^++Cl^-"

Let concentration of NCl is 0.1 M

Then concentration of Na+ is 0.1 M

And consideration of Cl- is 0.1 M


"Na_2SO_4 \rightarrow 2Na^++SO_4^{2-}"

If concentration of "Na_2SO_4" is 0.2 M,

Then concentration of Na+ is 0.2x2=0.4 M

And consideration of "SO_4^{2-}" is 0.2 M


If we have 1 l of this solution total concentration of Na+ is 0.1+0.4=0.5M

"I=0.5\sum c(z)^2=0.5(0.5(1)^2+0.1(-1)^2+0.2(-2)^2)=0.7 M"